The Ka of propanoic acid (C2H5COOH) is 1.34 × 10-5. Calculate the pH of the solution AND the concentrations of C2H5COOH AND C2H5COO– in a 0.323 M propanoic acid solution at equilibrium. Please explain each step in detail.
Answer #1


                  C2H5COOH ---> C2H5COO– + H+

initial            0.323          0         0

at equilibrium      0.323-x          x       x

kA = [C2H5COO– ][H+]/[C2H5COOH]

1.34 * 10^(-5) = x^2/(0.323-x)

x = 0.0021 M

[C2H5COO– ]= 0.021 M [H+] = 0.0021 M

[C2H5COOH] = 0.323-0.0021 = 0.3209 M